Radius of first orbit of hydrogen atom is 0.53 .Radius of fifth state of hydrogen aton is  
a 2.65
b 13.25
c 20
d infinite

Grade:12

2 Answers

sama
3 years ago

r(n)= n^2 * r(1)

r(5)= 5^2 * r(1) = 25*0.53 = 13.25 (b)
Radius of fifth state of hydrogen aton is 13.25 
Pawan Prajapati
3 years ago
Answer ::

Formula for Bohr's radius,
Rn = n^2 × R1

Here n = 5

R5 = 5^2 × 0.53

R5 = 13.25

Option B is correct.


Thanks